OpenStudy (anonymous):

What is a correct construction of a line parallel to given line l and passing through given point Q? (1 point)

OpenStudy (anonymous):

Do you have to only use a compass and ruler?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Put your compass on Q first and draw an arc so that it hits the line in two places

OpenStudy (anonymous):

At each of those intersecting points draw an arc through point q and make it large enough where the arcs also intersect on the other side

OpenStudy (anonymous):

connect the points where the arcs intersect, one of the points is Q. You have now created a perpendicular line

OpenStudy (anonymous):

now start on Q again. Draw an arc to that i intersects the new perpendicular line at 2 points.

OpenStudy (anonymous):

at each of these intersecting points draw a circle of the same size. Connect the points where these circles intersect

OpenStudy (anonymous):

You have now created another perpendicular line to the new line consequestly a parallel line to the original

OpenStudy (anonymous):

I did it exactly how you said and I got it right thanks! Can you help me with another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, 1) with slope –4. (1 point)

OpenStudy (anonymous):

Point slope form: y – y1 = m(x – x1) where m is the slope an (x1, y1)

OpenStudy (anonymous):

you tell me the answer

OpenStudy (anonymous):

y-1 = -4(x-4)??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok heres another one Write the equation for the horizontal line that contains point G(–8, 8).

OpenStudy (anonymous):

Horizontal line: y=c Vertical line: x=c where c is the constant coordinate

OpenStudy (anonymous):

y=8 and x=-8

OpenStudy (anonymous):

y=8 would be the horizontal

OpenStudy (anonymous):

so the final equation would just be y=8

OpenStudy (anonymous):

@ChmE